OpenStudy (anonymous):

how would you prove that 2 plus 2 equals 4

OpenStudy (zzr0ck3r):

the complete proof of 2 + 2 = 4 involves 2,452 subtheorems

OpenStudy (zzr0ck3r):

"One of the reasons that the proof of 2 + 2 = 4 is so long is that 2 and 4 are complex numbers—i.e. we are really proving (2+0i) + (2+0i) = (4+0i)—and these have a complicated construction (see the Axioms for Complex Numbers) but provide the most flexibility for the arithmetic in our set.mm database. In terms of textbook pages, the construction formalizes perhaps 70 pages from Takeuti and Zaring's detailed set theory book "

OpenStudy (zzr0ck3r):

still want the proof?

OpenStudy (anonymous):

oh my /the 1st. 327 pages of principia mathematica by sir bertrand russell/whitehead are devoted to proving 1 plus 1 equals 2 and they do do you have a link i go to or do you have the proof i found a very shrewd rather concise proof using composition of functions via set theoretic notation

OpenStudy (anonymous):

do you have the proof or a link i go to

OpenStudy (zzr0ck3r):

just google.